# Extension principles for the Einstein Yang–Mills system

### v1

@@ 4.1.2. Weighted energy estimates
$\int_{C_{u}}\xi^{1-\gamma}\frac{\Omega^{2}h}{r^{1+b}}\left(\Phi^{2}+(9+2b)\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}\underline{u}$
$+\int_{\underline{C}_{\underline{u}}}\xi^{1-\gamma}\frac{h^{-1}}{r^{1+b}}\left(2\left(w\alpha-\Phi\right)^{2}+\frac{\Phi^{2}}{2}+2h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u$
$+\iint_{\mathcal{S}}(\gamma-b)\xi^{1-\gamma}\frac{\Omega^{2}}{r^{2+b}}(\alpha^{\prime})^{2}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
$+\iint_{\mathcal{S}}\varepsilon\Lambda\xi^{1-\gamma}\frac{\Omega^{2}}{r^{2+b}}(\alpha^{\prime})^{2}\mathrm{d}u\mathrm{d}\underline{u}+\iint_{\mathcal{S}}\varepsilon\Lambda\xi^{1-\gamma}\frac{\Omega^{2}}{r^{2+b}}\left(\alpha^{2}+\Phi^{2}\right)\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}\xi^{-\gamma}\frac{\Omega^{2}}{r^{2+b}}\left(\xi\alpha^{2}+\left(w\alpha-\Phi\right)^{2}+\Phi^{2}+2h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u\mathrm{d}\underline{u}.$
Due to the choice $\gamma>b$, the second term on the right hand side having the small constant $\varepsilon\Lambda$ can be absorbed by the spacetime integral on the left hand side. To bound the other spacetime integrals on the right hand side, we appeal to the first order estimates (4.1.2) and (4.1.2) so that we are able to deduce the energy bound
$\int_{C_{u}}\xi^{1-\gamma}\frac{\Omega^{2}h}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}\underline{u}$
$+\int_{\underline{C}_{\underline{u}}}\xi^{1-\gamma}\frac{h^{-1}}{r^{1+b}}\left(\left(w\alpha-\Phi\right)^{2}+\Phi^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u$
$+\iint_{\mathcal{S}}\frac{(\gamma-b)}{2}\xi^{1-\gamma}\frac{\Omega^{2}}{r^{2+b}}(\alpha^{\prime})^{2}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
Once the energy estimates are completed, we are able to improve the pointwise estimates.
4.1.3. Pointwise estimates
In the following lemmas 4.54.7, we will improve the $L^{\infty}$ estimates for $r^{-1}\hat{w},\,\alpha$, $\mu$ and $\big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\big{|}$ to close the bootstrap argument.
Lemma 4.5.
We have the refine estimate
$\xi^{-1-\gamma}\mu\lesssim\varepsilon^{2}r^{b},\quad\text{in}\,\,\mathcal{S}.$
Proof.
The proof is analogous to that of Lemma 3.4. By the mass formula,
$\frac{m}{r^{1+b}\xi^{1+\gamma}}={}$ $\frac{1}{r^{1+b}\xi^{1+\gamma}}\int_{r_{\ast}}^{r}\left((1-\mu)\alpha^{2}+\Phi^{2}\right)\mathrm{d}r^{\prime}|_{C_{u}},$
[…]
where in the second line we have used the fact $\partial_{r}\xi\geq 0$ along $C_{u}\cap\mathcal{S}$. ∎
Lemma 4.6.
We have in $\mathcal{S}$
$\xi^{-\gamma}\alpha^{2}\lesssim\varepsilon^{2}r^{b},\quad\xi^{-2-\gamma}r^{-2}\hat{w}^{2}\lesssim\varepsilon^{2}r^{b}.$
Proof.
We will first show that $\|\alpha\|^{2}_{H^{1}_{w}(C_{u_{0}})}\lesssim\varepsilon^{2}$ implies
$\|\xi^{-\gamma}r^{-b}\alpha^{2}\|_{L^{\infty}(C_{u_{0}}\cap\mathcal{S})}\lesssim\varepsilon^{2}.$ (4.9)
Integrating along $\partial_{r}$ on $C_{u_{0}}$, we obtain the following identity
$\xi^{-\gamma}r^{-b}\alpha^{2}\big{|}_{C_{u_{0}}}(\underline{u})-\xi^{-\gamma}r^{-b}\alpha^{2}\big{|}_{C_{u_{0}}}(\underline{u}_{\ast})$
[…]
Noting that $\xi^{-\gamma}r^{-b}\alpha^{2}\big{|}_{C_{u_{0}}}(\underline{u}_{\ast})=0$, we have
$\xi^{-\gamma}r^{-b}\alpha^{2}\big{|}_{C_{u_{0}}}(\underline{u})+\int_{C_{u_{0}}}\gamma\frac{\alpha^{2}}{\xi^{1+\gamma}r^{1+b}}\mathrm{d}r$
[…]
which justifies (4.9).
Next, we prove pointwise estimate for $\alpha$, using (4.9).
Multiplying $2\xi^{-\gamma}r^{-b}\alpha$ on (2.15) and integrating along $\partial_{u}$ on each $\underline{C}_{\underline{u}}$, noting that $\gamma>b$, we have
$\big{|}\xi^{-\gamma}r^{-b}\alpha^{2}-\xi^{-\gamma}r^{-b}\alpha^{2}|_{u_{0}}\big{|}+\int_{\underline{C}_{\underline{u}}}\left(\gamma-b\right)(1-\mu)\xi^{-\gamma}\frac{\alpha^{2}}{r^{1+b}}h^{-1}\mathrm{d}u$
[…]
$+\int_{\underline{C}_{\underline{u}}}\gamma\Big{|}r\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\Big{|}\xi^{-\gamma}\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}u$
[…]
$+\int_{\underline{C}_{\underline{u}}}\left(2\mu+\gamma\big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\big{|}\right)\xi^{-\gamma}\frac{\alpha^{2}}{r^{1+b}}h^{-1}\mathrm{d}u.$
By the bootstrap assumptions, $\big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\big{|}$ and $\mu$ are both small enough, then
$(\gamma-b)(1-\mu)>\mu+\gamma\big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\big{|}.$
After absorbing the error terms on the right hand side, we conclude
$\xi^{-\gamma}r^{-b}\alpha^{2}+\int_{\underline{C}_{\underline{u}}}\frac{\gamma-b}{2}\xi^{-\gamma}\frac{\alpha^{2}}{r^{1+b}}h^{-1}\mathrm{d}u$
[…]
where in the last inequality we have appealed to (4.9) and the energy estimates (4.1.2) and (4.1.2).
Meanwhile, integrating $\xi^{-2-\gamma}r^{-b-2}\hat{w}^{2}$ along $\partial_{\underline{u}}$ on each $C_{u}$ gives rise to
$\xi^{-2-\gamma}r^{-b-2}\hat{w}^{2}{}$ $+(2+\gamma)\int_{C_{u}}\xi^{-3-\gamma}(1-\xi)\frac{r^{-2}\hat{w}^{2}}{r^{1+b}}\Omega^{2}h\mathrm{d}\underline{u}$
$+(2+b)\int_{C_{u}}\xi^{-2-\gamma}\frac{r^{-2}\hat{w}^{2}}{r^{1+b}}\Omega^{2}h\mathrm{d}\underline{u}$
$\lesssim$ $\int_{C_{u}}\left(\xi^{-1-\gamma}\frac{\alpha^{2}}{r^{1+b}}+\xi^{-3-\gamma}\frac{r^{-2}\hat{w}^{2}}{r^{1+b}}\right)\Omega^{2}h\mathrm{d}\underline{u}\lesssim\varepsilon^{2}.$
We complete the proof of this lemma. ∎
Lemma 4.7.
We have the improvement,
$\Big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\Big{|}\lesssim\varepsilon^{2},\quad\text{in}\,\,\mathcal{S}.$
Proof.
By means of a straightforward calculation, using the mean value theorem, we find that
$rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}={}$ $\frac{rh\underline{h}-r_{\ast}h_{\ast}\underline{h}_{\ast}+r_{\ast}h_{\ast}\underline{h}_{\ast}-rh\underline{h}_{\ast}}{r-r_{\ast}}$
[…]
which is small in view of Lemma 4.6 and the fact
$\big{|}\ln\frac{h|_{r_{2}}}{h_{\ast}}\big{|}\lesssim\int^{r_{2}}_{r_{\ast}}\frac{\alpha^{2}}{r}\mathrm{d}r\lesssim\varepsilon^{2}.$
What is more, along the lines of the subsection 3.1.6, we are able to prove the boundedness of the following geometric quantities,
$|\ln(\Omega^{2}h)|+|\ln(-\underline{h})|+|\ln h|.$
In the end, we present the pointwise estimates for $\partial_{u}w$ and $\alpha^{\prime}$.
Lemma 4.8.
In the case of $0<b<\frac{\sqrt{89}-9}{4}$, there is
$\xi^{-3-\gamma}|\partial_{u}w|^{2}\lesssim\varepsilon^{2}r^{b}.$
The proof resembles the one in the second part of Lemma 4.6.
Lemma 4.9.
Suppose $\xi^{2-\gamma}r^{-b}(\alpha^{\prime})^{2}\big{|}_{C_{u_{0}}}$ is bounded, then we have the following $L^{\infty}$ estimate for $\alpha^{\prime}$,
$\xi^{2-\gamma}(\alpha^{\prime})^{2}\lesssim r^{b}.$
Proof.
Multiply $2\xi^{2-\gamma}r^{-b}\alpha^{\prime}$ on the second order equation (3.1.4) and integrate along $\partial_{u}$. Then
$\xi^{2-\gamma}r^{-b}(\alpha^{\prime})^{2}+\int_{\underline{C}_{\underline{u}}}(\gamma-b)\xi^{2-\gamma}\frac{(\alpha^{\prime})^{2}}{r^{1+b}}h^{-1}\mathrm{d}u$
[…]
$+\int_{\underline{C}_{\underline{u}}}\left(c+\varepsilon\right)\xi^{2-\gamma}\frac{(\alpha^{\prime})^{2}}{r^{1+b}}h^{-1}\mathrm{d}u,$
where we take the constant $c>0$ to be small, so that $c+\varepsilon$ can be absorbed by the positive integral on the left hand side, noting $02\geq\gamma>b>$ 0. Thus this lemma is concluded.
4.2. Cauchy stability in weighted $H^{1}$ space
The proof of Theorem 4.1 itself leads to the Cauchy stability theorem in $H^{1}_{w}$ space.
Theorem 4.10.
[…]
$\int_{C^{[\underline{u}_{\ast},\underline{u}_{1}]}_{u_{0}}}\frac{\Omega^{2}h}{r^{1+b}}\left(\frac{r^{-2}\hat{w}^{2}}{\xi^{3+\gamma}}+\frac{\alpha^{2}}{\xi^{1+\gamma}}+\xi^{1-\gamma}\frac{(\alpha^{\prime})^{2}}{2}\right)\mathrm{d}\underline{u}<\varepsilon^{2},$
and assume $|r-r_{\ast}|^{-\frac{\gamma}{2}}\alpha|_{C_{u_{0}}}$ is continuous at $\underline{u}=\underline{u}_{\ast}$. There exists an $\varepsilon_{0}>0$ such that if $\varepsilon^{2}\leq\varepsilon^{2}_{0}$, the following conclusions hold.
Case I: $b>0$. There is a unique solution
$(r,\,m,\,\alpha)\in C^{1}\times C^{1}\times H^{1}_{w}\quad\text{in}\,\,\mathcal{D}(u_{0},\,u;\,\underline{u}_{1}),\,\,u\in[u_{0},\,\underline{u}_{\ast}).$
Moreover, for any $(u,\,\underline{u})\in[u_{0},\,\underline{u}_{\ast})\times[\underline{u}_{\ast},\,\underline{u}_{1}]$, there is
$\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(\frac{(\partial_{u}w)^{2}}{\xi^{3+\gamma}}+\frac{r^{-2}\hat{w}^{2}}{\xi^{1+\gamma}}+\xi^{1-\gamma}\alpha^{2}\right)\mathrm{d}u$
$+\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}\left(\frac{r^{-2}\hat{w}^{2}}{\xi^{3+\gamma}}+\frac{\alpha^{2}}{\xi^{1+\gamma}}+\xi^{1-\gamma}(\alpha^{\prime})^{2}\right)\mathrm{d}\underline{u}\lesssim\varepsilon^{2},$
and we have the following pointwise estimates
$\xi^{-2-\gamma}r^{-2}\hat{w}^{2}+\xi^{-\gamma}\alpha^{2}+\xi^{-3-\gamma}|\partial_{u}w|^{2}+\xi^{-1-\gamma}\mu\lesssim\varepsilon^{2}r^{b}.$
Case II: $b=0$. There is a unique solution
$(r,\,m,\,\alpha)\in C^{1}\times C^{0}\times H^{1}_{w},\quad\text{in}\,\,\mathcal{D}(u_{0},\,u;\,\underline{u}_{1}),\,\,u\in[u_{0},\,\underline{u}_{\ast}).$
The estimates in Case I hold (with $b=0$) except the $L^{\infty}$ estimate for $\partial_{u}w$.
Appendix A Local existence theorem
For the characteristic problem where initial data are prescribed on a future light cone with vertex at the center of symmetry, the local existence theorem for the classical solution to the spherically symmetric Einstein scalar field system was proved in [8]. Regarding to the proof, Christodoulou reduced the whole system in Bondi coordinates to a scalar equation and utilized the contracting mapping principle to obtain the existence theorem. However, in this section, we mean to provide an alternative proof based on the energy method. To be more precisely, we work within the double null gauge and demonstrate a local existence theory for the SSEYM system in $H^{1}$ (and $H^{1}_{w}$) space, using an iteration method.
Theorem A.1.
Let $\underline{u}_{1}\in(u_{0},\,+\infty)$ be a finite constant. Suppose on the initial slice $C_{u_{0}}$, the following energy is bounded
$\int_{C_{u_{0}}}r^{-1-b}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r<I_{1}.$
Case I: $0<b<\frac{\sqrt{89}-9}{4}$. i) There exists $u_{\ast}>u_{0}$ such that the SSEYM system has a unique solution $\left(r,\,m,\,\alpha\right)\in C^{1}\times C^{1}\times H^{1}$ on the domain $\mathcal{D}(u_{0},u_{\ast};\underline{u}_{1})$, which enjoys the subsequent bound,
$\int_{C_{u}}r^{-1-b}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r\lesssim I_{1},$
$\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+h^{2}(\partial_{u}\hat{w})^{2}\right)\mathrm{d}u\lesssim I_{1},$
$|r^{-1}\hat{w}|^{2}+|\partial_{u}\hat{w}|+|\alpha|^{2}+\mu\lesssim I_{1}r^{b},$
ii) The solution depends continuously on data in the following sense. Suppose $\left(r_{n},\,m_{n},\,\hat{w}_{n}\right)$ and $\left(r_{m},\,m_{m},\,\hat{w}_{m}\right)$ are solutions with data $\hat{w}_{0n}$, $\hat{w}_{0m}$ respectively, then
$\int_{C_{u}}\left(\frac{|\hat{w}_{n}-\hat{w}_{m}|^{2}}{r^{3+b}_{n}}+\frac{|\alpha_{n}-\alpha_{m}|^{2}}{r^{1+b}_{n}}+\frac{|\alpha^{\prime}_{n}-\alpha^{\prime}_{m}|^{2}}{r^{1+b}_{n}}\right)\mathrm{d}r_{n}$
$+\int_{\underline{C}_{\underline{u}}}\frac{h_{n}^{-1}}{r^{1+b}_{n}}\left(\frac{\left(\hat{w}_{n}-\hat{w}_{m}\right)^{2}}{r_{n}^{2}}+\left(\alpha_{n}-\alpha_{m}\right)^{2}+h_{n}^{2}(\partial_{u}\hat{w}_{n}-\partial_{u}\hat{w}_{m})^{2}\right)\mathrm{d}u$
[…]
Case II: $b=0$. The statement in Case I holds with $b=0$, except that the pointwise estimate for $|\partial_{u}w|$ is absent and hence in this case the solution $\left(r,\,m,\,\alpha\right)\in C^{1}\times C^{0}\times H^{1}$.
Remark A.2.
We can prove the local existence theorem in $H^{1}_{w}$ space (or rather the space of data in Theorem 4.1) with $0\leq b<\frac{\sqrt{89}-9}{4}$ and $b<\gamma\leq 2$ straightforwardly in the same manner as Theorem A.1. Thus, we omit the details.
Alternatively, since $H^{1}$ is dense in $H^{1}_{w}$ (actually, smooth functions with compact support are dense in $H^{1}_{w}$, see Appendix C), we can make use of the property ii) that the solution continuously depends on data, to verify the local existence in $H^{1}_{w}$ space.
Proof.
For the sake of simplicity, we only carry out the proof for the case of $b=0$ in details. The proof for the other case $0<b<\frac{\sqrt{89}-9}{4}$ goes through in an analogous fashion.
It suffices to prove the local existence theorem near the axis, which essentially turns the local existence problem into the one with small data. Away from the axis, the local existence follows in a standard way.
Choose $\underline{u}_{1}>u_{0}$ and $\underline{u}_{1}-u_{0}$ sufficiently small so that the initial energy is sufficiently small. We denote the size of data by $\varepsilon^{2}I_{1}$, where $\varepsilon^{2}$ is small enough, namely,
$\int_{C_{u_{0}}^{[u_{0},\,\underline{u}_{1}]}}r^{-1-b}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r\leq\varepsilon^{2}I_{1}.$
We shall find a local solution for the SSEYM system using the method of Picard iteration. Notice that, the Yang–Mills equation can be alternatively written in terms of $\hat{w}$ (1.3)
$\partial_{u}\partial_{\underline{u}}\hat{w}+2\frac{\Omega^{2}}{r^{2}}\hat{w}=\Omega^{2}\frac{\hat{w}^{2}}{r^{2}}\left(\hat{w}+3\right).$ (A.1)
[…]
$\partial_{u}\partial_{\underline{u}}\hat{w}_{k+1}+{}$ $2\Omega_{k}^{2}\frac{\hat{w}_{k+1}}{r_{k}^{2}}=\Omega_{k}^{2}\frac{\hat{w}_{k}^{2}}{r_{k}^{2}}\left(\hat{w}_{k}+3\right),$ (A.2)
$\partial_{\underline{u}}m_{k+1}={}$ $\Omega^{2}_{k}h_{k}\left(\frac{\Phi_{k+1}^{2}}{2}-h_{k}\underline{h}_{k}\alpha_{k+1}^{2}\right),$ (A.3)
$\partial_{u}m_{k+1}={}$ $\underline{h}_{k}\left(\frac{\Phi_{k+1}^{2}}{2}-\left(h_{k}\underline{h}_{k}\right)^{-1}h^{2}_{k}(\partial_{u}\hat{w}_{k+1})^{2}\right),$ (A.4)
$\partial_{u}\left(\Omega^{2}_{k+1}h_{k+1}\right)={}$ $\Omega^{2}_{k+1}\underline{h}^{-1}_{k+1}\underline{h}_{k}\left(-\frac{2m_{k+1}}{r^{2}_{k}}+\frac{\Phi_{k+1}^{2}}{r_{k}}\right),$ (A.5)
$\partial_{\underline{u}}\left(\underline{h}_{k+1}\right)={}$ $\Omega^{2}_{k}h_{k}h_{k+1}^{-1}\left(-\frac{2m_{k+1}}{r^{2}_{k}}+\frac{\Phi_{k+1}^{2}}{r_{k}}\right),$ (A.6)
with initial data prescribed in Theorem A.1 and the boundary conditions
@@ Appendix A Local existence theorem
In the linearized system (A.2)-(A.6), we first use (A.2) to solve $\hat{w}_{k+1}$. Next, substituting the $\hat{w}_{k+1}$ into (A.3) and (A.4), we can further solve $m_{k+1}$. After that, substituting both of $\hat{w}_{k+1}$ and $m_{k+1}$ into (A.5) and (A.6), we are able to solve $\Omega^{2}_{k+1}h_{k+1}$ and $\underline{h}_{k+1}$, and hence obtain $r_{k+1}$.
Auxiliary equations. Along with the main equations, we can derive the following two equations from (A.3) and (A.6), (A.4) and (A.5).
$\partial_{\underline{u}}h_{k+1}={}$ $2\Omega^{2}_{k}h_{k}(1-\mu_{k})\underline{h}_{k+1}^{-1}\frac{(\alpha_{k+1})^{2}}{r_{k}},$ (A.8)
$\partial_{u}(\Omega^{-2}_{k+1}\underline{h}_{k+1})={}$ $-2\Omega^{-2}_{k+1}h^{-1}_{k+1}h_{k}\frac{(\partial_{u}\hat{w}_{k+1})^{2}}{r_{k}}.$ (A.9)
@@
Uniform bound. We claim that if $\varepsilon$ (or rather $\underline{u}_{1}-u_{0}$) is small enough, there is a constant $M<+\infty$ such that we have the uniform bounds, for all $i\geq 0$,
@@
[…]
When $i=0$, an energy argument (in line with the proof leading to (3.1.4) in the subsection 3.1.4) for the linearized equation
@@
[…]
Thus, (A) holds for $i=0$, which further implies
@@
To proceed, let us suppose the estimate (A) holds for $i\leq k-1$ with the constant $M$ to be determined later. Then we can derive the pointwise estimates, as in the subsections 3.1.5 and 3.1.6
@@
Moreover, using the equation of (A.5) with $k$ replaced by $k-1$ or $k-2$, and noting the gauge choice (A.7), we have
@@
In addition, taking the equation of (A.6) (with $k$ replaced by $k-1$) and the gauge condition of $\underline{h}_{k}|_{\Gamma}=-\Omega^{2}_{k}h_{k}|_{\Gamma}$ into account, we find that
@@
We shall then prove (A) holds for $i=k$. Thanks to the preceding estimates and an energy argument in line with the subsection 3.1.4, we are able to derive
@@
[…]
where the universal constant $C$ is independent of $k$. If we choose $M>CI_{1}$, then (A) with $i=k$ is verified.
Convergence. In the compact domain concerned, since $\{r_{k}\}$ is equi-continuous, there is a subsequence, which we still denote by $\{r_{k}\}$, such that $r_{k}\rightarrow r$ on the compact neighbourhood. In what follows, we always use this converging subsequence to replace the original one.
Based on (A.2), we derive the following linear equation,
@@
[…]
@@
$\|U_{k}-U_{k-1}\|_{L^{\infty}}\leq M2^{-k+1},\quad\|\mu_{k}-\mu_{k-1}\|_{L^{\infty}}\leq M2^{-k+1},$ (A.13)
$\Big{|}\int_{C_{u}}\frac{m_{k}}{r_{k-1}^{2}}\mathrm{d}r_{k-1}|_{C_{u}}-\int_{C_{u}}\frac{m_{k-1}}{r_{k-2}^{2}}\mathrm{d}r_{k-2}|_{C_{u}}\Big{|}\leq M2^{-k+1},$ (A.14)
$\Big{|}\int_{\underline{C}_{\underline{u}}}\frac{m_{k}}{r_{k-1}^{2}}\mathrm{d}r_{k-1}|_{\underline{C}_{\underline{u}}}-\int_{\underline{C}_{\underline{u}}}\frac{m_{k-1}}{r_{k-2}^{2}}\mathrm{d}r_{k-2}|_{\underline{C}_{\underline{u}}}\Big{|}\leq M2^{-k+1},$ (A.15)
for a large constant $M$ to be determined. When $k=1$, (A.13)-(A.15) hold by the uniform boundedness in the previous step. We will prove the same inequalities hold with $k$ replaced by $k+1$.
As is obvious from the expression, the linear equation (A) shares the same structure as (A.2). Thus, we can establish an analogous energy identity, noting that data vanishes now
@@
[…]
where, for simplicity, we denote
@@
[…]
@@
[…]
@@
[…]
For notational convenience, in what follows, we will denote

### v2

@@ 4. Local existence theorem
$\int_{C_{u_{0}}^{[u_{0},\,\underline{u}_{1}]}}r^{-1}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r\leq\varepsilon^{2}I_{1}.$
We shall find a local solution for the SSEYM system using the method of Picard iteration.
Linearized system. We let $\left(r_{0},\,\hat{w}_{0},\,m_{0}\right)=\left(r|_{C_{u_{0}}}+\frac{u_{0}-u}{2},\,0,\,0\right)$ and then define $\left(r_{k+1},\,m_{k+1},\,\hat{w}_{k+1}\right)$ iteratively for $k\geq 0$ as the unique solution to the linearized system:
$\partial_{u}\partial_{\underline{u}}\hat{w}_{k+1}+{}$ $2\Omega_{k}^{2}\frac{\hat{w}_{k+1}}{r_{k}^{2}}=\Omega_{k}^{2}\frac{\hat{w}_{k}^{2}}{r_{k}^{2}}\left(\hat{w}_{k}+3\right),$ (4.2)
$\partial_{\underline{u}}m_{k+1}={}$ $\Omega^{2}_{k}h_{k}\left(\frac{\Phi_{k+1}^{2}}{2}-h_{k}\underline{h}_{k}\alpha_{k+1}^{2}\right),$ (4.3)
$\partial_{u}m_{k+1}={}$ $\underline{h}_{k}\left(\frac{\Phi_{k+1}^{2}}{2}-\left(h_{k}\underline{h}_{k}\right)^{-1}h^{2}_{k}(\partial_{u}\hat{w}_{k+1})^{2}\right),$ (4.4)
$\partial_{u}\left(\Omega^{2}_{k+1}h_{k+1}\right)={}$ $\Omega^{2}_{k+1}\underline{h}^{-1}_{k+1}\underline{h}_{k}\left(-\frac{2m_{k+1}}{r^{2}_{k}}+\frac{\Phi_{k+1}^{2}}{r_{k}}\right),$ (4.5)
$\partial_{\underline{u}}\left(\underline{h}_{k+1}\right)={}$ $\Omega^{2}_{k}h_{k}h_{k+1}^{-1}\left(-\frac{2m_{k+1}}{r^{2}_{k}}+\frac{\Phi_{k+1}^{2}}{r_{k}}\right),$ (4.6)
with initial data prescribed in Theorem 4.1 and subject to the boundary conditions
@@
In the linearized system (4.2)-(4.6), we first solve (4.2) to determine $\hat{w}_{k+1}$. Substituting the $\hat{w}_{k+1}$ into (4.3) and (4.4), we then solve $m_{k+1}$. Finally, substituting both of $\hat{w}_{k+1}$ and $m_{k+1}$ into (4.5) and (4.6), we are able to solve $\Omega^{2}_{k+1}h_{k+1}$ and $\underline{h}_{k+1}$, thereby obtaining $r_{k+1}$.
Auxiliary equations. Along with the main equations, we can derive the following two equations from (4.3) and (4.6), (4.4) and (4.5).
$\partial_{\underline{u}}h_{k+1}={}$ $2\Omega^{2}_{k}h_{k}(1-\mu_{k})\underline{h}_{k+1}^{-1}\frac{(\alpha_{k+1})^{2}}{r_{k}},$ (4.8)
$\partial_{u}(\Omega^{-2}_{k+1}\underline{h}_{k+1})={}$ $-2\Omega^{-2}_{k+1}h^{-1}_{k+1}h_{k}\frac{(\partial_{u}\hat{w}_{k+1})^{2}}{r_{k}}.$ (4.9)
@@
Uniform bound. We assert that if $\varepsilon$ (or equivalently, $\underline{u}_{1}-u_{0}$) is small enough, there is a constant $M<+\infty$ such that the uniform bounds,
@@
$\leq{}$ $\varepsilon^{2}M,\quad u\in[u_{0},\,u_{\ast}],\,\,\underline{u}\in[u,\,\underline{u}_{1}],$ (4.10)
hold for all $i\geq 0$,
When $i=0$, an energy argument (in line with the proof leading to (3.5) in the subsection 3.5) for the linearized equation
@@
$={}$ $\int_{C_{u_{0}}}\frac{\Omega^{2}_{0}h_{0}}{r_{0}}\left((\alpha^{\prime}_{1})^{2}+9\alpha_{1}^{2}+4r_{0}^{-2}\hat{w}_{1}^{2}\right)\mathrm{d}\underline{u}.$
Thus, (4) holds for $i=0$, which further implies
@@
To proceed, let us suppose the estimate (4) holds for $i\leq k-1$ with the constant $M$ to be determined later. Then we can derive the pointwise estimates, as in the subsections 3.6 and 3.7
@@
Moreover, using the equation of (4.5) with $k$ replaced by k1k-1 or k2k-2, and noting the gauge choice (4.7), we have
@@
In addition, taking the equation of (4.6) (with $k$ replaced by k1k-1) and the gauge condition of $\underline{h}_{k}|_{\Gamma}=-\Omega^{2}_{k}h_{k}|_{\Gamma}$ into account, we find that
@@
We shall then prove (4) holds for $i=k$. Thanks to the preceding estimates and an energy argument analogous to that in Subsection 3.5, we are able to derive
@@
$\leq{}$ $\varepsilon^{2}I_{1}\cdot C,$ (4.11)
where the universal constant $C$ is independent of $k$. If we choose $M>CI_{1}$, then (4) with $i=k$ is verified.
Convergence. In the compact domain concerned, the sequence $\{r_{k}\}$ is equicontinuous. Therefore, there is a subsequence (still denoted by $\{r_{k}\}$) that converges uniformly to a function $r$, $r_{k}\rightarrow r$, on the compact neighbourhood. In the subsequent analysis, we will always work with this convergent subsequence in place of the original sequence.
Based on (4.2), we derive the following linear equation,
@@
$={}$ $f\left(\frac{\hat{w}_{k+1}}{r_{k}},\,\frac{\hat{w}_{k}}{r_{k}},\,\frac{\hat{w}_{k-1}}{r_{k-1}}\right)\cdot\frac{(U_{k}-U_{k-1})}{r_{k}},$ (4.12)
@@
$\|U_{k}-U_{k-1}\|_{L^{\infty}}\leq M2^{-k+1},\quad\|\mu_{k}-\mu_{k-1}\|_{L^{\infty}}\leq M2^{-k+1},$ (4.13)
$\Big{|}\int_{C_{u}}\frac{m_{k}}{r_{k-1}^{2}}\mathrm{d}r_{k-1}|_{C_{u}}-\int_{C_{u}}\frac{m_{k-1}}{r_{k-2}^{2}}\mathrm{d}r_{k-2}|_{C_{u}}\Big{|}\leq M2^{-k+1},$ (4.14)
$\Big{|}\int_{\underline{C}_{\underline{u}}}\frac{m_{k}}{r_{k-1}^{2}}\mathrm{d}r_{k-1}|_{\underline{C}_{\underline{u}}}-\int_{\underline{C}_{\underline{u}}}\frac{m_{k-1}}{r_{k-2}^{2}}\mathrm{d}r_{k-2}|_{\underline{C}_{\underline{u}}}\Big{|}\leq M2^{-k+1},$ (4.15)
for a large constant $M$ to be determined. When $k=1$, (4.13)-(4.15) hold by the uniform boundedness in the previous step. We will prove the same inequalities hold with $k$ replaced by $k+1$.
As is evident from its expression, the linear equation (4) shares the same structure as (4.2). Thus, we can establish an analogous energy identity, taking into account that the initial data now vanish,
@@
$\lesssim{}$ $\iint_{\mathcal{D}}\left(\big{|}\frac{\hat{w}_{k+1}}{r_{k}}\big{|}+\big{|}\frac{\hat{w}_{k}}{r_{k-1}}\big{|}+\big{|}\frac{\hat{w}_{k-1}}{r_{k-2}}\big{|}\right)\Omega_{k-1}^{2}\frac{|U_{k}-U_{k-1}||V_{k+1}-V_{k}|}{r_{k-1}r_{k-2}}\mathrm{d}u\mathrm{d}\underline{u}.$
Here, for simplicity, we denote
@@
$\lesssim{}$ $\|U_{k}-U_{k-1}\|_{L^{\infty}}\left(\iint_{\mathcal{D}}\frac{\Omega_{k-1}^{2}}{r^{2}_{k-2}}\left(\big{|}\frac{\hat{w}_{k+1}}{r_{k}}\big{|}^{2}+\big{|}\frac{\hat{w}_{k}}{r_{k-1}}\big{|}^{2}+\big{|}\frac{\hat{w}_{k-1}}{r_{k-2}}\big{|}^{2}\right)\mathrm{d}u\mathrm{d}\underline{u}\right)^{\frac{1}{2}}$
@@
$\lesssim{}$ $\varepsilon\|U_{k}-U_{k-1}\|_{L^{\infty}}\left(\iint_{\mathcal{D}}\Omega_{k-1}^{2}\frac{|V_{k+1}-V_{k}|^{2}}{r^{2}_{k-1}}\mathrm{d}u\mathrm{d}\underline{u}\right)^{\frac{1}{2}}.$
@@
$\lesssim{}$ $\varepsilon\|U_{k}-U_{k-1}\|^{2}_{L^{\infty}}.$ (4.16)
For notational simplification, in what follows, we will denote